# XAI-CF – Examining the Role of Explainable Artificial Intelligence in Cyber Forensics

### v2

@@ Abstract
With the rise of complex cyber devices such as smartphones, Internet of Things, and control systems in automotive and drones, the proliferation of operating systems and file formats, pervasive encryption, use of the cloud for remote processing and storage, and legal standards, Cyber Forensics (CF) is facing many new challenges. For example, there are dozens of systems running on smartphones, each with more than millions of downloadable applications. Sifting through this large amount of data and making sense requires new techniques, such as from the field of Artificial Intelligence (AI). To apply these techniques successfully in CF, we need to justify and explain the results to the stakeholders of CF, such as forensic analysts and members of the court, for them to make an informed decision. If we want to apply AI successfully in CF, there is a need to develop trust in AI systems. Some other factors in accepting the use of AI in CF are to make AI authentic, interpretable, understandable, and interactive. This way, AI systems will be more acceptable to the public and ensure alignment with legal standards. An explainable AI (XAI) system can play this role in CF, and we call such a system XAI-CF. XAI-CF is indispensable and is still in its infancy. In this paper, we explore and make a case for the significance and advantages of XAI-CF. We strongly emphasize the need to build a successful and practical XAI-CF system and discuss some of the main requirements and prerequisites of such a system. We present a formal definition of the terms CF and XAI-CF and a comprehensive literature review of previous works that apply and utilize XAI to build and increase trust in CF. To make the reader familiar with the study carried out in this paper, in addition to background, we also present a critical and short review of the works carried out in the last ten years in XAI and CF. We discuss some challenges facing XAI-CF, such as adversarial attacks, bias management, oversimplification, CF and AI chasm, and human-computer interaction. We also provide some concrete solutions to these challenges. We identify key insights and future research directions for building XAI applications for CF. This paper is an effort to explore and familiarize the readers with the role of XAI applications in CF, and we believe that our work provides a promising basis for future researchers interested in XAI-CF.
@@
Artificial intelligence, Cyber forensics, Explainable artificial intelligence, Trustworthy cyber forensics, XAI-CF.
journal: Forensic Science International: Digital Investigation
@@ 1 Introduction and Motivation
Forensics analysis is a major part of investigating a security incident and involves detecting, collecting, and documenting the pieces of evidence. Most of the time forensics analysis is linked with evidence to the court of law. When these pieces of evidence are found on a cyber device then their analysis is called Cyber Forensics (CF). One of the first documented and practical examples of CF investigation was carried out in 1986 (stoll1988stalking, ; stoll2005cuckoo, ). The intruder was caught by tracing the network activities and performing a forensics analysis of different computer logs. At that time forensics was performed by computer professionals on an ad-hoc case-by-case basis. The forensics analyst in the case (stoll1988stalking, ) was an Astronomer (stoll2005cuckoo, ).
@@
With the rise of complex cyber devices such as smartphones, Internet of Things, and control systems in automotive and drones, the proliferation of operating systems and file formats, pervasive encryption, use of the cloud for remote processing and storage, and legal standards, CF is facing many new challenges. For example, there are dozens of systems running on smartphones, each with more than millions of downloadable applications. Sifting through this large amount of data and making sense requires new techniques, such as from the field of Artificial Intelligence (AI). A detailed list and discussion of challenges in CF can be found in (garfinkel2010digital, ; du2020sok, ).
@@
The first documented research proposal on AI (AI-modern-approach-2020, ) was initiated by McCarthy et al. (mccarthy2006proposal, ) in 1955. The major goals discussed in the proposal were: automation; how to program a computer to use a language; neuron nets for forming concepts; and self-improvement. Some of the current traditional goals of AI research include reasoning; decision-making; knowledge representation; natural language processing; and Machine Learning (ML). Some of the major techniques that AI uses to achieve these goals are statistics; artificial neural networks; formal logic; cognitive psychology; and linguistics. Most of the recent breakthroughs in modern AI came through deep learning (AI-modern-approach-2020, ). AI is a powerful technology and has potential benefits but also comes with potential risks and challenges, such as lack of transparency, inherently unexplainable especially when using deep learning, adversarial attacks, privacy, misinformation, algorithm and data bias, and ethics. For example, the recently developed ChatGPT (ChatGPT, ) the new AI language model has revolutionized the approach in AI and is being used in various domains, such as business, education, healthcare, etc. The potentials of ChatGPT are unlimited, but it can be a double-edged sword (fui2023generative, ). For example, for businesses, it can create creative content such as ideas for advertisements, and on the other hand, it can create hallucinations and produce fake information. Another example is the use of ChatGPT in higher education which can promote collaboration among students and remote learning, and can also be used for improving student assessments, such as grading essays, assignments, exams, and quizzes more quickly and accurately (cotton2023chatting, ). However, using ChatGPT for assessment comes with some challenges, such as the possibility of plagiarism (cotton2023chatting, ). ChatGPT can generate essays based on a set of parameters or prompts, which means students can generate and submit essays that are not their own work (dehouche2021plagiarism, ). It becomes very difficult to distinguish between a student’s own writing and the responses generated by ChatGPT. Moreover, the responses generated by ChatGPT may not accurately reflect the true level of understanding of the student. For a detailed list of risks and challenges, interested readers are referred to (scherer2015regulating, ; AI-modern-approach-2020, ; fui2023generative, ; ray2023chatgpt, ).
@@
AI has been successfully applied in various fields as well as in Cyber Security, but its application and utilization specifically in CF is still relatively limited (automation-ai-df-2021, ). To bridge this gap we need to examine and explore the potential effects of AI in CF and develop standardized techniques for mining cyber (digital) evidence (df-ai-2022, ). Due to the complexities involved in the process of CF, which requires an intelligent analysis sifting through and making sense of different patterns, the use of AI in this field is particularly valuable (ai-df-2014, ). The lack of transparency of an AI system gives rise to different challenges in the field of CF. One of the main challenges is that it requires validation by a human user (hall2022explainable, ). This opacity makes it difficult to admit a CF analysis using AI in a court of law because it is often taken as a black box that is impossible to justify or explain up to legal standards (ai-df-2014, ).
@@
One of the most important factors in accepting the use of AI in CF is developing trust in AI systems. Some of the other factors in accepting the use of AI in CF are, to make AI authentic, interpretable, understandable, and interactive. This way, AI systems will be more acceptable to the public and ensure alignment with legal standards. An explainable AI (XAI) system can play this role in CF, and we call such a system XAI-CF. XAI adds transparency and accountability to an AI system and helps create trustworthy AI. This helps a human investigator to understand and validate AI systems/algorithms, which makes them more reliable and trustworthy (xai-cybersecurity-survey-2022, ). XAI rationalizes the workings and justifies the outputs of an AI system (hall2022explainable, ). XAI-CF is indispensable and is still in its infancy. In this paper, we explore and make a case for the significance and advantages of XAI-CF. We strongly emphasize the need to build a successful and practical XAI-CF system and discuss some of the main requirements and prerequisites of such a system.
@@
There are different stakeholders of XAI (preece2018stakeholders, ). Each one of them has different requirements for XAI. Developers of AI applications, want XAI to aid in testing, debugging, and evaluation. Theorists who are advancing AI theory, want to understand XAI to advance the state of the art in XAI rather than develop practical XAI applications. Ethicists who are policy makers, and critics of AI, are concerned about the accountability and transparency of XAI. Users of XAI are the most influential stakeholders and contribute the most to growing the literature on XAI. In terms of the CF, they are part of the users’ stakeholders. Members of the CF community need explanations to better understand the output of an AI system decide how to act on the given outputs and justify their actions. These actions play an important role in making key and accountable decisions in a court of law.
@@
The graphs shown in Figure 1 depict the popularity by month of the two terms Cyber Forensics and Explainable Artificial Intelligence when used during Google searches performed in the last ten years (2014 – 2023). This shows the development in the use of terminology based on the current most popular search engine that is useful and of value to identify trends. Starting from the year 2014 the term Cyber Forensics remains popular and even gains in popularity after the year 2021. There are few mentions of the term Explainable Artificial Intelligence in the years from 2014 – 2020. After 2020 the term XAI starts gaining popularity. Still, XAI is not considered a popular term in comparison to CF, and there is a need to develop awareness in the general public of its importance and relevance in specific domains. This is one of the motivations of this paper to make a case for the significance and advantages of XAI in CF and strongly recommend its use in CF. Based on the popularity results from this graph the study in this paper reviews the works on XAI, CF, and XAI-CF only for the last ten years (2014 – 2023).
@@
Figure 1: Google search trends (Google-search-trends, ) by month for the last ten years, for the terms Explainable Artificial Intelligence and Cyber Forensics. The y-axis depicts the relative search frequency for the term. A value of 100 is the peak popularity for the term. A value of 50 means that the term is half as popular. A score of 0 means there was not enough data for this term.
@@ 1.1 Contributions
The following are the major contributions of this paper:
@@
    To make the reader familiar with the study carried out in this paper, in addition to background, we also present a critical and short review of the works carried out in the last ten years in XAI and CF.
@@
    We discuss some challenges facing XAI-CF, such as adversarial attacks, bias management, oversimplification, CF and AI chasm, and human-computer interaction, and also provide some concrete solutions to these challenges.
@@ 1.2 Definitions
Before defining CF, we present standard definitions of the terms Cyber and Forensics by the Cambridge Dictionary. Cyber (def-cyber, ) – Computers, especially the Internet, and activities that use them. Forensics (def-forensics, ) – Scientific methods of solving crimes, that involve examining objects or substances related to a crime. Based on these definitions, we define:
@@
[…]
@@
Four main groups form the major Stakeholders of CF: (1) developers of CF; (2) CF analysts; (3) members of the court, such as the judges, lawyers, and jury members; (4) and anyone legally involved in the case, such as the accused, perpetrators, and victims. We have included the fourth group as the stakeholders because they are also interested in learning how the AI system works. This can help them carry out adversarial attacks using anti-forensics techniques [76] to evade detection by the XAI system during forensics analysis.
@@ 1.3 Organization of the Paper
The rest of the paper is organized as follows. Section 2 describes the methods used to carry out the survey in this paper. Section 3 details the related works (works similar to our work), i.e., the survey/review of the works about the application of XAI in CF. Section 4 give background information about XAI and CF and also presents a short critical review of last ten year of XAI and CF. Section 5 discusses why, how, evaluation, and desiderata of XAI-CF. Section 6 presents a comprehensive literature review of previous works carried out in the area of XAI-CF. Section 7 discusses the challenges of building a practical XAI-CF system/framework and provides solutions to these challenges as part of future works. Section 8 discusses the threats to the validity of the study presented in this paper. Finally, section 9 concludes our work.
@@ 2 Research Methodology
The goal of this work is to examine and investigate the research carried out in the area of XAI-CF. Figure 2 shows the flowchart of the research methodology used for the survey carried out in this paper. The flowchart is developed and used to carry out the four surveys/reviews in this paper. Two short reviews, one on the Last Ten Years of XAI in section 4.2 and second on the Last Ten Years of CF in section 4.4. Two comprehensive reviews, one is the Related Work in section 3 and the second is the Literature Review in section 6. To increase the credibility of the study carried out in this paper, we used the principle of triangulation (yin2015qualitative, ; patton2014qualitative, ), i.e., seeking at least three ways of verifying a procedure, piece of data, or finding.
@@
Then, we examine whether the Publication Year of each selected article falls within the range of 2014 to 2023. If the answer is yes, indicating the relevance and recency of the literature, the flowchart proceeds to the next step, where the articles undergo an English language check. However, if the answer is no, the flowchart directs the articles to the Eliminate stage, as they may not align with the desired time frame for the research. This step ensures that the chosen articles are not only pertinent to the research focus but also recent enough to reflect the current state of XAI in the context of CF. After the English language comes the Duplicate Check stage. If the article is duplicated it is directed to the eliminate stage. After this stage, the total number of articles selected for classification and review were 66. Distribution of these selected articles by source and year is shown in Figure 3. These articles are reviewed in sections 3 , 4.2 , 4.4 , and 6 in the paper.
[…]
Figure 3: Distribution of the $6$ selected articles for classification and review by source and year. The source Others include one paper (1.52%) each from Wiley, AAAI, PeerJ, NDSS, Hindawi, Intechopen, IOS, PMLR, ISM, and ISMIR.
@@ 3 Related Work
There are several works (adadi2018peeking, ; arrieta2020explainable, ; javed2023survey, ) that have reviewed/surveyed the research works carried out on the application of XAI in different domains and fields. In this section, we discuss the previous works that have specifically explored (surveyed/reviewed) in some context the application and utilization of XAI in CF.
@@
The work carried out in (kelly2020explainable, ) studies the challenges of developing XAI systems to automate the decision-making process in the generation of evidence in CF. The challenges discussed, such as a large number of features, missing data, etc., are mostly related to XAI systems. However, One of the challenges of how to drive a decision tree automatically using the available forensic evidence data is specifically related to CF. There is no comprehensive survey or review, of previous and current works, presented in the paper about the application and utilization of XAI in CF. The case study of drug testing presented in the paper collects medical forensics lab data, such as hair and nails, through a structured questionnaire, and then digitizes it for forensic analysis.
@@
The work carried out in (solanke2022explainable, ) proposes recommendations for an interpretable AI-based CF. Some of the recommendations for mitigating distrust in AI discussed in the paper are: as far as possible use less complex, non-opaque techniques; the methods should be evaluated based on their ability to articulate the explanation; and incorporation of a human-in-the-loop so that validations can be performed before reaching a decision. The paper also highlights some of the methods for explaining AI models, such as model simplification, feature importance, visualization, etc. Like the previous work (kelly2020explainable, ) this paper highlights the shortcomings and challenges of AI and how they can be improved for successfully applying AI in digital forensics. There is no comprehensive survey on the application of XAI in CF.
@@
Recently several works (rjoub2023survey, ; xai-cybersecurity-survey-2022, ; zhang2022explainable, ; charmet2022explainable, ) present a comprehensive survey on the application and utilization of XAI in Cyber Security. The main focus of the works is on Cyber Security, and the only work that briefly discusses CF is (xai-cybersecurity-survey-2022, ). All of these works present a detailed literature review of the works that seek to achieve explainability in the following sub-fields of Cyber Security: Intrusion detection systems; Malware detection; Phishing and Spam detection; and BotNet detection. (xai-cybersecurity-survey-2022, ) presents a short discussion about fraud detection, zero-day vulnerabilities, digital forensics, cyber-physical systems, and crypto-jacking.
@@
The work carried out in (hall2022explainable, ) is the closest to our work. They also explore the potential of XAI to enhance CF. The main focus of this study is to present a model of XAI that can be applied to a CF case for investigative support. The authors also present a fictional case study of XAI for investigative support. The paper also discusses the future of AI as XAI in CF and how it is going to change the field of CF. However, there is no comprehensive survey or review, of previous and current works, presented in the paper about the application and utilization of XAI in CF.
@@
Table 1 presents a comparison of previous surveys with the survey carried out in this paper. As we can see, unlike other works, the work carried out in this paper covers all the major aspects of a successful and practical XAI-CF system. If we want to apply AI successfully in CF then we need to develop trust in AI systems by making them authentic, interpretable, understandable, and interactive. The role of XAI in CF is indispensable and we need more research work in this area. This is the main motivation of this paper to explore the role and application of XAI-CF. Based on our exhaustive search and discussion presented above, this is the first work that carries out a comprehensive survey/review of the application and utilization of XAI-CF.
@@
Table 1: Comparison of our survey/review with seven other similar works discussed in Section 3
Survey Year
XAI in CF Detailed Survey
XAI in CF Challenges
XAI in CF How to Improve
Main Contributions
Our Survey 2024
Yes
Yes
Yes
@@
short survey XAI; short survey CF; detailed literature review of XAI in CF; desiderata, challenges and future of XAI in CF;
Rjoub et al. (rjoub2023survey, ) 2023
No
No
No
detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
Hall et al. (hall2022explainable, ) 2022
No
Yes
Yes
model of XAI for investigative support; future of XAI in CF; case study;
Capuano et al. (xai-cybersecurity-survey-2022, ) 2022
No
Yes
No
@@
short discussion on XAI in CF; detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
Zhang et al. (zhang2022explainable, ) 2022
No
No
No
detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
Charmet et al. (charmet2022explainable, ) 2022
No
No
No
detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
Solanke (solanke2022explainable, ) 2022
No
Yes
Yes
@@
Highlights methods and tools to explain AI models for use in CF; proposed structure of XAI model to be used in CF;
Kelly et al. (kelly2020explainable, ) 2020
No
Yes
No
@@
short survey of XAI in CF; XAI in automating decision tree in CF; challenges of XAI in CF; case study;

### v3

@@ Abstract
With the rise of complex cyber devices, Cyber Forensics (CF) is facing numerous new challenges. For instance, many systems are operating on smartphones, each hosting millions of downloadable applications. Analyzing this vast amount of data and making sense of it requires innovative techniques, particularly those from Artificial Intelligence (AI). To successfully implement these techniques in CF, it is essential to justify and explain the results to CF stakeholders. This will enable informed decision-making and foster trust in AI systems. An explainable AI (XAI) system can fulfill this role in CF, and we refer to this system as XAI-CF. Although XAI-CF is crucial, it is still in its early stages of development. Thus, there is a need to investigate and advocate for the importance and benefits of XAI-CF. Furthermore, previous studies do not provide a comprehensive survey of XAI’s role in CF. Many only briefly address the challenges and ways to enhance XAI for this field, and none have proposed a unified framework for an XAI-CF system. To address these research gaps, this paper discusses the key requirements and prerequisites for an XAI-CF system. We present a thorough literature review of prior works that apply and utilize XAI to build and enhance trust in CF. In addition to discussing the challenges faced by XAI-CF and offering concrete solutions, we introduce the first XAI-CF framework that cohesively integrates XAI principles across every stage of the CF lifecycle. We believe that our work provides a solid foundation for future researchers interested in XAI-CF.
@@
Artificial intelligence, Cyber forensics, Explainable artificial intelligence, Trustworthy cyber forensics, Explainable artificial intelligence in cyber forensics.
@@ 1 Introduction and Motivation
Forensic analysis is a major part of investigating a security incident and involves detecting, collecting, and documenting the pieces of evidence. Most of the time, forensic analysis is linked with evidence for the court of law. When these pieces of evidence are found on a cyber device, then their analysis is called Cyber Forensics (CF). One of the first documented and practical examples of CF investigation was carried out in 1986 [1, 2]. The intruder was caught by tracing the network activities and performing a forensics analysis of different computer logs. At that time, forensics was performed by computer professionals on an ad-hoc case-by-case basis. The forensics analyst in the case [1] was an Astronomer [2].
@@
With the rise of complex cyber devices such as smartphones, Internet of Things, and control systems in automotive and drones, the proliferation of operating systems and file formats, pervasive encryption, use of the cloud for remote processing and storage, and legal standards, CF is facing many new challenges. For example, there are dozens of systems running on smartphones, each with more than a million downloadable applications. Sifting through this large amount of data and making sense requires new techniques, such as those from the field of Artificial Intelligence (AI). A detailed list and discussion of challenges in CF can be found in [3, 4].
@@
The first documented research proposal on AI [5] was initiated by McCarthy et al. [6] in 1955. The major goals discussed in the proposal were: automation; how to program a computer to use a language; neuron nets for forming concepts; and self-improvement. Some of the current traditional goals of AI research include reasoning; decision-making; knowledge representation; natural language processing; and Machine Learning (ML). Some of the major techniques that AI uses to achieve these goals are statistics; artificial neural networks; formal logic; cognitive psychology; and linguistics. Most of the recent breakthroughs in modern AI came through deep learning [5]. AI is a powerful technology and has potential benefits, but also comes with potential risks and challenges, such as lack of transparency, inherently unexplainable, especially when using deep learning, adversarial attacks, privacy, misinformation, algorithm and data bias, and ethics. For example, the recently developed ChatGPT [7], the new AI language model, has revolutionized the approach in AI and is being used in various domains, such as business, education, healthcare, etc. The potentials of ChatGPT are unlimited, but it can be a double-edged sword [8]. For example, for businesses, it can create creative content such as ideas for advertisements, and on the other hand, it can create hallucinations and produce fake information. Another example is the use of ChatGPT in higher education which can promote collaboration among students and remote learning, and can also be used for improving student assessments, such as grading essays, assignments, exams, and quizzes more quickly and accurately [9]. However, using ChatGPT for assessment comes with some challenges, such as the possibility of plagiarism [9]. ChatGPT can generate essays based on a set of parameters or prompts, which means students can generate and submit essays that are not their own work [10]. It becomes very difficult to distinguish between a student’s own writing and the responses generated by ChatGPT. Moreover, the responses generated by ChatGPT may not accurately reflect the true level of understanding of the student. For a detailed list of risks and challenges, interested readers are referred to [11, 5, 8, 12].
@@
AI has been successfully applied in various fields as well as in Cyber Security, but its application and utilization specifically in CF is still relatively limited [13]. To bridge this gap, we need to examine and explore the potential effects of AI in CF and develop standardized techniques for mining cyber (digital) evidence [14]. Due to the complexities involved in the process of CF, which requires an intelligent analysis sifting through and making sense of different patterns, the use of AI in this field is particularly valuable [15]. The lack of transparency of an AI system gives rise to different challenges in the field of CF. One of the main challenges is that it requires validation by a human user [16]. This opacity makes it difficult to admit a CF analysis using AI in a court of law because it is often taken as a black box that is impossible to justify or explain up to legal standards [15].
@@
One of the most important factors in accepting the use of AI in CF is developing trust in AI systems. Some of the other factors in accepting the use of AI in CF are to make AI authentic, interpretable, understandable, and interactive. This way, AI systems will be more acceptable to the public and ensure alignment with legal standards. An explainable AI (XAI) system can play this role in CF, and we call such a system XAI-CF. XAI adds transparency and accountability to an AI system and helps create trustworthy AI. This helps a human investigator to understand and validate AI systems/algorithms, which makes them more reliable and trustworthy [17]. XAI rationalizes the workings and justifies the outputs of an AI system [16]. XAI-CF is indispensable and is still in its infancy. In this paper, we explore and make a case for the significance and advantages of XAI-CF. We strongly emphasize the need to build a successful and practical XAI-CF system and discuss some of the main requirements and prerequisites of such a system.
@@
There are different stakeholders of XAI [18]. Each one of them has different requirements for XAI. Developers of AI applications want XAI to aid in testing, debugging, and evaluation. Theorists who are advancing AI theory want to understand XAI to advance the state of the art in XAI rather than develop practical XAI applications. Ethicists who are policy makers, and critics of AI, are concerned about the accountability and transparency of XAI. Users of XAI are the most influential stakeholders and contribute the most to growing the literature on XAI. In terms of the CF, they are part of the users’ stakeholders. Members of the CF community need explanations to better understand the output of an AI system, decide how to act on the given outputs, and justify their actions. These actions play an important role in making key and accountable decisions in a court of law.
@@
The graphs shown in Figure 1 depict the popularity by month of the two terms Cyber Forensics and Explainable Artificial Intelligence when used during Google searches performed in the last ten years (2016 – 2025). This shows the development in the use of terminology based on the current most popular search engine, which is useful and of value to identify trends. Starting from the year 2016 the term Cyber Forensics remains popular and even gains in popularity after the year 2022. There are a few mentions of the term Explainable Artificial Intelligence in the years from 2016 – 2021. After 2021, the term XAI started gaining popularity. In 2025, XAI has seen a substantial rise in popularity, although its search volume is still roughly 50% lower than that of CF. Still, XAI is not considered a popular term in comparison to CF, and there is a need to develop awareness in the general public of its importance and relevance in specific domains. This is one of the motivations of this paper to make a case for the significance and advantages of XAI in CF and strongly recommend its use in CF. Based on the popularity results from this graph, the study in this paper reviews the works on XAI, CF, and XAI-CF only for the last ten years (2016 – 2025).
@@
Figure 1: Google search trends [19] by month for the last ten years, for the terms Explainable Artificial Intelligence and Cyber Forensics. The y-axis depicts the relative search frequency for the term. A value of 100 is the peak popularity for the term. A value of 50 means that the term is half as popular. A score of 0 means there was not enough data for this term.
@@ 1.1 Contributions
Despite the rising interest in XAI from 2021 to 2025, no survey has focused solely on its application in CF. Existing works do not provide a comprehensive review of XAI’s role in cyber forensics, and only a handful briefly mention the challenges and potential improvements for applying XAI in this area. In this paper, we examine and justify the importance and benefits of XAI-CF. This work concentrates specifically on XAI in the context of CF. Readers seeking broader discussions on the role and impact of AI in CF may consult [20, 21, 22]. The main contributions of this paper are as follows:
@@
    To make the reader familiar with the study carried out in this paper, in addition to the background, we also present a critical and short review of the works carried out in the last ten years in XAI and CF.
@@
    We discuss some challenges facing XAI-CF, such as adversarial attacks, bias management, oversimplification, the CF and AI chasm, and human-computer interaction, and also provide some concrete solutions to these challenges.
@@
    The paper presents the first XAI-CF framework that systematically integrates XAI principles throughout all phases of the cyber forensics lifecycle.
  • 6.
@@ 1.2 Definitions
Before defining CF, we present standard definitions of the terms Cyber and Forensics by the Cambridge Dictionary. Cyber [23] – Computers, especially the Internet, and activities that use them. Forensics [24] – Scientific methods of solving crimes, that involve examining objects or substances related to a crime. Based on these definitions, we define:
@@
Before defining XAI-CF in this paper, we present the definition of XAI from other sources. The first definition presented by David Gunning [25] states “The Explainable AI (XAI) program aims to create a suite of ML techniques that enable human users to understand, appropriately trust, and effectively manage the emerging generation of artificially intelligent partners.” Another definition is by Arrieta et al. [26] that states “Given an audience, an XAI system produces details or reasons to make its functioning clear or easy to understand.” The term Explainable means able to explain, and the Cambridge Dictionary defines the term Explain [27] – To make something clear or easy to understand by describing or giving information about it. Based on all these definitions, we define:
@@
Four main groups form the major Stakeholders of CF: (1) developers of CF; (2) CF analysts; (3) members of the court, such as the judges, lawyers, and jury members; (4) and anyone legally involved in the case, such as the accused, perpetrators, and victims. We have included the fourth group as stakeholders because they are also interested in learning how the AI system works. This can help them carry out adversarial attacks using anti-forensics techniques [76] to evade detection by the XAI system during forensics analysis.
@@ 1.3 Organization of the Paper
The rest of the paper is organized as follows. Section 2 describes the methods used to carry out the survey in this paper. Section 3 details the related works (works similar to our work), i.e., the survey/review of the works about the application of XAI in CF. Section 4 gives background information about XAI and CF and also presents a short critical review of the last ten years of XAI and CF. Section 5 discusses why, how, evaluation, and desiderata of XAI-CF. Section 6 presents a comprehensive literature review of previous works carried out in the area of XAI-CF. Section 7 discusses the challenges of building a practical XAI-CF system/framework and provides solutions to these challenges as part of future work. Section 8 discusses the threats to the validity of the study presented in this paper. Finally, section 9 concludes our work. Figure 2 provides an overview of the paper’s structure and progression.
Figure 2: Overview of the paper’s structure and progression.
@@ 2 Research Methodology
The goal of this work is to examine and investigate the research carried out in the area of XAI-CF. Figure 3 shows the flowchart of the research methodology used for the survey carried out in this paper. The flowchart is developed and used to carry out the four surveys/reviews in this paper. Two short reviews, one on the Last Ten Years of XAI in section 4.2 and the second on the Last Ten Years of CF in section 4.4. Two comprehensive reviews, one is the Related Work in section 3 and the second is the Literature Review in section 6. To increase the credibility of the study carried out in this paper, we used the principle of triangulation [28, 29], i.e., seeking at least three ways of verifying a procedure, piece of data, or finding.
@@
Then, we examine whether the Publication Year of each selected article falls within the range of 2016 to 2025. If the answer is yes, indicating the relevance and recency of the literature, the flowchart proceeds to the next step, where the articles undergo an English language check. However, if the answer is no, the flowchart directs the articles to the Eliminate stage, as they may not align with the desired time frame for the research. This step ensures that the chosen articles are not only pertinent to the research focus but also recent enough to reflect the current state of XAI in the context of CF. After the English language comes the Duplicate Check stage. If the article is duplicated, it is directed to the elimination stage. After this stage, the total number of articles selected for classification and review was 69. Distribution of these selected articles by source and year is shown in Figure 4. These articles are reviewed in sections 3 , 4.2 , 4.4 , and 6 in the paper.
Elsevier27.5%IEEE27.5%Springer10.2%MDPI8.7%ArXiv5.8%Nature5.8%Others 14.5%
20167.2%20174.4%20185.8%20195.8%202017.4%20217.2%202220.3%202311.6%202410.2%2025 10.2%
Figure 4: Distribution of the 69 selected articles for classification and review by source and year. The source Others include one paper (1.5%) each from ACM, OUP, Wiley, AAAI, PeerJ, NDSS, Hindawi, IntechOpen, PLOS, and ISMIR.
@@ 3 Related Work
There are several works [30, 26, 31] that have reviewed/surveyed the research works carried out on the application of XAI in different domains and fields. In this section, we discuss the previous works that have specifically explored (surveyed/reviewed) in some context the application and utilization of XAI in CF.
@@
The work carried out in [32] studies the challenges of developing XAI systems to automate the decision-making process in the generation of evidence in CF. The challenges discussed, such as a large number of features, missing data, etc., are mostly related to XAI systems. However, one of the challenges of how to drive a decision tree automatically using the available forensic evidence data is specifically related to CF. There is no comprehensive survey or review of previous and current works presented in the paper about the application and utilization of XAI in CF. The case study of drug testing presented in the paper collects medical forensics lab data, such as hair and nails, through a structured questionnaire, and then digitizes it for forensic analysis.
@@
The work carried out in [33] proposes recommendations for an interpretable AI-based CF. Some of the recommendations for mitigating distrust in AI discussed in the paper are: as far as possible, use less complex, non-opaque techniques; the methods should be evaluated based on their ability to articulate the explanation; and the incorporation of a human-in-the-loop so that validations can be performed before reaching a decision. The paper also highlights some of the methods for explaining AI models, such as model simplification, feature importance, visualization etc. Like the previous work [32] this paper highlights the shortcomings and challenges of AI and how they can be improved for successfully applying AI in digital forensics. There is no comprehensive survey on the application of XAI in CF.
@@
Recently, several works [34, 35, 36, 37, 38, 17, 39, 40] present a comprehensive survey on the application and utilization of XAI in Cyber Security. The main focus of the works is on Cyber Security, and the only work that briefly discusses CF is [17]. All of these works present a detailed literature review of the works that seek to achieve explainability in the following sub-fields of Cyber Security: Network security, Endpoint security; Application security, Intrusion detection systems, Malware detection, Phishing and Spam detection, and BotNet detection. [17] presents a short discussion about fraud detection, zero-day vulnerabilities, digital forensics, cyber-physical systems, and crypto-jacking.
@@
The work carried out in [16] is the closest to our work. They also explore the potential of XAI to enhance CF. The main focus of this study is to present a model of XAI that can be applied to a CF case for investigative support. The authors also present a fictional case study of XAI for investigative support. The paper also discusses the future of AI as XAI in CF and how it is going to change the field of CF. However, there is no comprehensive survey or review of previous and current works presented in the paper about the application and utilization of XAI in CF.
@@
Table 1 presents a comparison of previous surveys with the survey carried out in this paper. As we can see, unlike other works, the work carried out in this paper covers all the major aspects of a successful and practical XAI-CF system. If we want to apply AI successfully in CF, then we need to develop trust in AI systems by making them authentic, interpretable, understandable, and interactive. The role of XAI in CF is indispensable, and we need more research work in this area. This is the main motivation of this paper to explore the role and application of XAI-CF. Based on our extensive search and the discussion provided above, this is the first study to deliver a comprehensive survey and review of XAI-CF applications and usage, while also proposing the first unified XAI-CF framework that spans all stages of the cyber forensics process.
@@
Table 1: Comparison of our survey/review with seven other similar works discussed in Section 3.
Survey Year XAI in CF Detailed Survey XAI in CF Challenges XAI in CF How to Improve Unified XAI-CF Framework Main Contributions
@@
Our Survey 2025 Yes Yes Yes Yes short survey XAI; short survey CF; detailed literature review of XAI in CF; desiderata, challenges and future of XAI in CF;
Rjoub et al. [38] 2023 No No No No detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
Hall et al. [16] 2022 No Yes Yes No model of XAI for investigative support; future of XAI in CF; case study;
@@
Capuano et al. [17] 2022 No Yes No No short discussion on XAI in CF; detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
Zhang et al. [39] 2022 No No No No detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
Charmet et al. [40] 2022 No No No No detailed survey of XAI in Cyber Security; future and challenges of XAI in Cyber Security;
@@
Solanke [33] 2022 No Yes Yes No highlights methods and tools to explain AI models for use in CF; proposed structure of XAI model to be used in CF;
@@
Kelly et al. [32] 2020 No Yes No No short survey of XAI in CF; XAI in automating decision tree in CF; challenges of XAI in CF; case study;